Find the slope of a line crossing (-1,2) and (-7,10). Then use it to write an equation for the line in slope- intercept form.
Answer by Cromlix(4381)

Hi there,
Slope (gradient) = y2 - y1/x2 -x1 = 10 -2/-7-(-1)) = 8/-6 = -4/3
Equation formula = y - b = m(x - a) You can use either of the coordinate sets. (-1, 2) or (-7, 10). I will use (-1, 2)
y - 2 = -4/3(x -(-1))= y - 2 = -4/3x -4/3 = y = -4/3x - 4/3 + 6/3
Answer = y = -4/3x + 2/3 or you can multiply throughout by 3
then answer = 3y = -4x + 2.
Hope this helps :-)